OpenStudy (anonymous):

If g(x) = .5x + 3, find the value of g(0) + 2. A. 5 B. 4 C. 5.5 D. 7

OpenStudy (anonymous):

g(0) = 3 g(0) + 2 = 3 + 2 + 5

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

help :(

OpenStudy (anonymous):

g(x) = .5x + 3 g(0) = .5 x 0 + 3 = 3 g(0) + 2 = 3+2 = 5

OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line that passes through (4, 2) with a slope of 3. A. y = 3x - 10 B. y = 3x + 4 C. y - 2 = 3 (x - 4) D. y = 4x + 3 ??